# Unveiling Topological Structures in Text: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v2

@@ Abstract
The surge of data available on the internet has led to the adoption of various computational methods to analyze and extract valuable insights from this wealth of information. Among these, the field of Machine Learning (ML) has thrived by leveraging data to extract meaningful insights. However, ML techniques face notable challenges when dealing with real-world data, often due to issues of imbalance, noise, insufficient labeling, and high dimensionality. To address these limitations, some researchers advocate for the adoption of Topological Data Analysis (TDA), a statistical approach that discerningly captures the intrinsic shape of data despite noise. Despite its potential, TDA has not gained as much traction within the Natural Language Processing (NLP) domain compared to structurally distinct areas like computer vision. Nevertheless, a dedicated community of researchers has been exploring the application of TDA in NLP, yielding 87 papers we comprehensively survey in this paper. Our findings categorize these efforts into theoretical and non-theoretical approaches. Theoretical approaches aim to explain linguistic phenomena from a topological viewpoint, while non-theoretical approaches merge TDA with ML features, utilizing diverse numerical representation techniques. We conclude by exploring the challenges and unresolved questions that persist in this niche field. Resources and a list of papers on this topic can be found at: https://github.com/AdaUchendu/AwesomeTDA4NLP.  [^1] DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Department of the Air Force under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Air Force. © 2024 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work.
@@ 1 Introduction
TDA is a “collection of powerful tools that can quantify shape and structure in data” [^2] https://www.indicative.com/resource/topological-data-analysis/ and is inspired by the algebra and geometry mathematical fields. The benefits of TDA are vast, including the ability to extract additional features, typically not captured by other feature extraction techniques Uchendu et al. (2024); Michel et al. (2017). We call these features that TDA extracts - topological features. Unsurprisingly, since TDA is used to capture topological features, it has been applied to many tasks where data has distinct graphical structures Papamarkou et al. (2024); Hensel et al. (2021). These include tasks that have obvious graph-like structures such as protein classification Dey and Mandal (2018); Lamine et al. (2023); Valeriani et al. (2024) and drug discovery Alagappan et al. (2016); to those that are not so obvious such as diabetes classification Wamil et al. (2023); Skaf and Laubenbacher (2022), image classification Horn et al. (2022); Trofimov et al. (2023), time series analysis Petri and Leitao (2020); Tymochko et al. (2021); Gholizadeh and Zadrozny (2018), and speech processing Tulchinskii et al. (2023). However, since the shape of a text is not apparent, it has not gained as much attention in Natural Language Processing (NLP) as it has done in the Computer Vision Horn et al. (2022); Trofimov et al. (2023) and Medical Singh et al. (2023); Nielson et al. (2015) domains. Still, several researchers have found ways to extract additional features using TDA that the other typical numerical representation techniques in text such as TF-IDF, Word2Vec embeddings, BERT embeddings, etc have not captured. Figure 1 shows that there has been a gradual acceleration in the number of published works in TDA for NLP applications, and we project this trend will continue in the future.
@@
Figure 1: Number of NLP papers using TDA published each year from 2013 to 2024
@@
TDA aims to answer the main question - what is the true shape of a data? We survey 87 papers that have attempted to find an answer through various approaches. We categorize all observed approaches that incorporate TDA in NLP into two categories, namely theoretical Karlgren et al. (2014); Port et al. (2018) and non-theoretical Zhu (2013); Doshi and Zadrozny (2018). Theoretical approaches involve using TDA to explain linguistic phenomena by probing the topological space, shape, and evolution of topics. On the other hand, Non-theoretical approaches mainly discuss how to effectively apply existing numerical representation techniques in NLP to extract novel topological features with TDA. We will first discuss the principles behind TDA and the two main techniques employed for TDA feature extraction: Persistent Homology and Mapper.
@@ 2 Topological Data Analysis (TDA)
Topology is defined as “the study of geometric properties and spatial relations unaffected by the continuous change of shape or size of figures,” (Oxford Dictionary). TDA is then a collection of powerful techniques that can quantify the shape and structure of data Munch (2017). Two main techniques are used to extract TDA features - Persistent Homology and Mapper.
@@ 2.1 Persistent Homology
Using the method described above, data is represented as a point cloud, and circles are drawn around each point. Next, the radius of each circle is increased using a defined range of points, such that if the circles get bigger and touch, one of the points disappears and this is recorded as a $death$. Additionally, this process of perturbation in different dimensions can cause the $birth$ of a new hole which is also recorded. Thus, due to these deformations, the following TDA features can be extracted and recorded in a 3-column matrix, which consists of columns representing - the $birth$ (formation of holes), $death$ (deformation or the closing of holes), and persistence features. Persistence is defined as the length of time it took a feature to disappear or die ($death-birth$). The $death$ is recorded with the radii value at which the points overlap. Lastly, TDA features are typically visualized in a persistence diagram which is a visual representation of the 3-column matrix of TDA features. Figure 2 illustrates an example of the process of extracting TDA features using Persistent Homology. Other ways of visualizing TDA features include Persistence images Adams et al. (2017) and Barcode plots Ghrist (2008). Application-wise, PH has been used to extract novel features to complement existing NLP representations and improve various text classification performances Doshi and Zadrozny (2018); Uchendu et al. (2024); Wu et al. (2022).
@@ 3.1.1 Semantic Topological Space
In addition, Cavaliere et al. (2017) extracts main concepts from the texts by probing the context-aware semantic topological space built with simplicial complexes. Finally, Wagner et al. (2012) uses TF-IDF to numerically represent the top 10-50 words in a corpus and build a topological space that analyzes the structure of similarities within several documents. This topological space is built using discrete Morse theory and persistent homology to find meaningful topological patterns Wagner et al. (2012). However, as of 2012, they found that their technique was unsuccessful due to the computational costs, which is a testament to how the NLP field has improved such that we now have more tractable solutions, such as dimensionality reduction algorithms McInnes et al. (2018), and TDA packages (Ripser Bauer (2021), Sklearn-TDA Saul and Tralie (2019), PHAT Bauer et al. (2017)), as well as compute resources to efficiently construct topological spaces from large or complex data.
@@ 3.1.2 Syntactic Topological Space
Port et al. (2018) analyzes how syntactic parameters are distributed over different language families, including Indo-European, Niger-Congo, Austronesian, and Afro-Asiatic families. Port et al. (2018) show that the three persistent connected components ($\beta_{0}$) in the Niger-Congo family represents its three subfamilies - Mande, Atlantic-Congo, and Kordofania. The syntactic topological structures of these languages also reveal the historical linguistic phenomena that the Hellenic branch played a role in the historical development of the Indo-European languages Port et al. (2018).
@@ 3.2 Topology of Topic Evolution
The topology of topic evolution refers to the study and representation of how topics, themes, or concepts develop and change over time within a given corpus of texts or discourse in a topological space/structure. This concept is particularly relevant in fields where understanding the temporal dynamics of topics can provide insights into trends, shifts in public opinion, or the development of scientific or cultural themes.
Sami and Farrahi (2017) utilizes TDA to visualize the relationship between words in a text block, words in a corpus, and text blocks in a corpus. This is done by visualizing both local context (i.e., each text block in a set of sentences), and global context (i.e., occurrence of extracted words in the corpus) features. These features are extracted by using circular topology to represent words. Sami and Farrahi (2017) also find that using the circular topology in 2D space, core words from the text block or corpus stay close to the center, and the explanatory words remain close to the circle’s periphery.
@@ 3.3 Topological “Shape” of Words
To compare the language phylogenies or evolutionary history of language in the Indo-European language family, Draganov and Skiena (2024) investigates the “shape” of words by using their embeddings as proxies. Using persistent homology, they construct language phylogenetic trees for over 81 Indo-European languages. Experiments reveal that: (1) the shape of the word embedding of a language carries historical and structural information, similar to Port et al. (2018, 2022)’s findings; (2) TDA methods can successfully capture aspects of the shape of language embeddings Draganov and Skiena (2024). Fitz (2022) introduces a novel terminology - word manifold, which is a simplicial complex, whose topological space captures grammatical structure expressed by the corpus. This is done by implementing a technique for generating topological structure directly from strings of words. Fitz (2022). Experiments reveal that the homotopy type of the word manifold is also influenced by linguistic structure Fitz (2022). Similarly, Fitz et al. (2024) measures the topological complexity (known as perforation) of hidden representation of Transformer-based models to understand their topological shapes.
Similar to Port et al. (2018, 2022), Dong (2024) extracts the topological shapes of languages. Specifically South American languages - the Nuclear-Macro-Jê (NMJ) and Quechuan families using TDA. By using techniques like multiple correspondence analysis (MCA) for dimension reduction of the categorical-valued dataset and persistent homology, Dong (2024) visualizes each language in the selected families as a point cloud. This forms the topological shape of the South American languages, such that languages close together are more similar. By comparing the topological shapes of the languages, it is observed that there are major distinctions between the Jê-proper and the non-Jê-proper languages, as well as the northern and southern Quechuan languages Dong (2024).
@@ 4.1 Tasks, Learning Types, Modality, and TDA techniques
We categorize the NLP tasks to which TDA has been applied into 7 categories: (1) Classification Zhu (2013); Tymochko et al. (2021); Tulchinskii et al. (2024); (2) Clustering and Topic Modeling Holmes (2020); Guan et al. (2016); (3) Sentiment and Semantic Analysis Michel et al. (2017); Rawson et al. (2022); Jakubowski et al. (2020); (4) Structure and Visualization Haghighatkhah et al. (2022); Rathore et al. (2023); (5) Health, Social, and Scholarly Analysis Effah (2017); Novak (2019); Petri and Leitao (2020); (6) Speech and Music Processing Bonafos et al. (2023); Sassone et al. (2022); and (7) Model Interpretation and Analysis Kostenok et al. (2023); Gourgoulias et al. (2024); Haim Meirom and Bobrowski (2022).
Recently TDA has been heavily applied to deepfake text detection Løvlie (2023); Tulchinskii et al. (2024); Kushnareva et al. (2024, 2021); Uchendu et al. (2024), linguistic/grammatical acceptability classification Cherniavskii et al. (2022); Proskurina et al. (2023); Jain et al. (2024); Perez and Reinauer (2022), word sense induction and disambiguation Rawson et al. (2022); Jakubowski et al. (2020); Temčinas (2018), and polysemy word classification Jakubowski et al. (2020); Triki (2021); Shehu (2024); Garcia (2022). All of these five subtasks are particularly interesting because the unique labels have very similar features, making them non-trivial to distinguish from each other, making TDA a promising toolset, resulting in comparable performance to or even outperforming baselines.
Finally, we can also categorize non-theoretical TDA applications in NLP by learning types, Modality, and TDA techniques. Learning types has supervised Elyasi and Moghadam (2019); Lavery et al. (2024), and unsupervised Spannaus et al. (2024); Bonafos et al. (2024); Modality has text Triki (2021); Kostenok et al. (2023), and Speech Sassone et al. (2022); and TDA techniques, has Persistent Homology Torres-Tramón et al. (2015); Cherniavskii et al. (2022), and Mapper Holmes (2020); Elyasi and Moghadam (2019). Due to the binary nature of these three categories, they are not very informative and thus unsuitable as the main taxonomy for the non-theoretical applications.

### v3

# Unveiling Topological Structures from Language: A Comprehensive Survey of Topological Data Analysis Applications in NLP

@@ Abstract
The surge of data available on the internet has led to the adoption of various computational methods to analyze and extract valuable insights from this wealth of information. Among these, the field of Machine Learning (ML) has thrived by leveraging data to extract meaningful insights. However, ML techniques face notable challenges when dealing with real-world data, often due to issues of imbalance, noise, insufficient labeling, and high dimensionality. To address these limitations, some researchers advocate for the adoption of Topological Data Analysis (TDA), a statistical approach that discerningly captures the intrinsic shape of data despite noise. Despite its potential, TDA has not gained as much traction within the Natural Language Processing (NLP) domain compared to structurally distinct areas like computer vision. Nevertheless, a dedicated community of researchers has been exploring the application of TDA in NLP, yielding 95 papers we comprehensively survey in this paper. Our findings categorize these efforts into theoretical and non-theoretical approaches. Theoretical approaches aim to explain linguistic phenomena from a topological viewpoint, while non-theoretical approaches merge TDA with ML features, utilizing diverse numerical representation techniques. We conclude by exploring the challenges and unresolved questions that persist in this niche field. Resources and a list of papers on this topic can be found at: https://github.com/AdaUchendu/AwesomeTDA4NLP [^1] DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Department of the Air Force under Air Force Contract No. FA8702-15-D-0001 or FA8702-25-D-B002. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Air Force. © 2025 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work..
@@ 1 Introduction
Alternative Method Description
Principal Component Analysis (PCA) Reduces dimensionality while preserving variance using orthogonal transformations.
t-Distributed Stochastic Neighbor Embedding (t-SNE) Projects high-dimensional data into 2D or 3D space while preserving local relationships.
Uniform Manifold Approximation and Projection (UMAP) Similar to t-SNE but faster and better at preserving global structure.
Multidimensional Scaling (MDS) Projects high-dimensional data into lower dimensions by preserving pairwise distances.
Autoencoders A type of Neural network that learns compressed representations of data via encoding-decoding processes.
Manifold Learning (Isomap, LLE, etc.) Captures intrinsic structures in high-dimensional data through graph-based techniques.
Graph-based Learning (Spectral Clustering, GNNs) Uses graphs to model relationships and structure within data.
Clustering Methods (DBSCAN, K-Means, HDBSCAN) Groups similar data points based on distance and density.
Kernel Methods (SVM, Kernel PCA) Uses non-linear mappings to extract complex structures in data.
Geometric Deep Learning Uses neural networks on non-Euclidean spaces like graphs and manifolds.
Geometric techniques (Delaunay triangulation, Convex hull) Uses geometrical techniques to extract characteristics of data by analyzing spacial boundaries
Table 1: Alternatives to Topological Data Analysis in NLP
@@
TDA is a “collection of powerful tools that can quantify shape and structure in data” [^2] https://www.indicative.com/resource/topological-data-analysis/ and is inspired by the algebra and geometry mathematical fields. The benefits of TDA are vast, including the ability to extract additional features, typically not captured by other feature extraction techniques Uchendu et al. (2024); Michel et al. (2017). We call these features that TDA extracts - topological features. Unsurprisingly, since TDA is used to capture topological features, it has been applied to many tasks where data has distinct graphical structures Papamarkou et al. (2024); Hensel et al. (2021). These include tasks that have obvious graph-like structures, such as protein classification Dey and Mandal (2018); Lamine et al. (2023); Valeriani et al. (2024) and drug discovery Alagappan et al. (2016); to those that are not so obvious such as diabetes classification Wamil et al. (2023); Skaf and Laubenbacher (2022), image classification Horn et al. (2022); Trofimov et al. (2023), time series analysis Petri and Leitao (2020); Tymochko et al. (2021); Gholizadeh and Zadrozny (2018), and speech processing Tulchinskii et al. (2023). However, since the shape of a text is not apparent, it has not gained as much attention in Natural Language Processing (NLP) as it has done in the Computer Vision Horn et al. (2022); Trofimov et al. (2023) and Medical Singh et al. (2023); Nielson et al. (2015) domains. Still, several researchers have found ways to extract additional features using TDA that the other typical numerical representation techniques in text such as TF-IDF, Word2Vec embeddings, BERT embeddings, etc have not captured. Figure 1 shows that there has been a gradual acceleration in the number of published works in TDA for NLP applications, and we project this trend will continue in the future.
@@
Figure 1: Number of NLP papers using TDA published each year from 2013 to May-2025
@@
TDA aims to answer the main question - what is the true shape of a data? We survey 95 papers that have attempted to find an answer through various approaches. We categorize all observed approaches that incorporate TDA in NLP into two categories, namely theoretical Karlgren et al. (2014); Port et al. (2018) and non-theoretical Zhu (2013); Doshi and Zadrozny (2018). Theoretical approaches involve using TDA to explain linguistic phenomena by probing the topological space, shape, and evolution of topics. On the other hand, Non-theoretical approaches mainly discuss how to effectively apply existing numerical representation techniques in NLP to extract novel topological features with TDA. We will first discuss the principles behind TDA and the two main techniques employed for TDA feature extraction: Persistent Homology and Mapper.
@@ 2 Topological Data Analysis (TDA)
Topology is defined as “the study of geometric properties and spatial relations unaffected by the continuous change of shape or size of figures,” (Oxford Dictionary). TDA is then a collection of powerful techniques that can quantify the shape and structure of data Munch (2017). While, there are alternatives to TDA (i.e., see Table 1), TDA is the only technique that can extract not only local but global features, preserve the shape and structure of the data, and is still effective even with insufficient data. Two main techniques are used to extract TDA features - Persistent Homology and Mapper.
@@ 2.1 Persistent Homology
Using the method described above, data is represented as a point cloud, and circles are drawn around each point. Next, the radius of each circle is increased using a defined range of points, such that if the circles get bigger and touch, one of the points disappears and this is recorded as a $death$. Additionally, this process of perturbation in different dimensions can cause the $birth$ of a new hole which is also recorded. Thus, due to these deformations, the following TDA features can be extracted and recorded in a 3-column matrix, which consists of columns representing - the $birth$ (formation of holes), $death$ (deformation or the closing of holes), and persistence features. Persistence is defined as the length of time it took a feature to disappear or die ($death-birth$). The $death$ is recorded with the radii value at which the points overlap. Lastly, TDA features are typically visualized in a persistence diagram which is a visual representation of the 3-column matrix of TDA features. Figure 2 illustrates an example of the process of extracting TDA features using Persistent Homology. Other ways of visualizing TDA features include Persistence images Adams et al. (2017) and Barcode plots Ghrist (2008). Application-wise, PH has been used to extract novel features to complement existing NLP representations and improve various classification performances Doshi and Zadrozny (2018); Uchendu et al. (2024); Wu et al. (2022).
@@ 3 Theoretical Approaches of TDA in NLP
Theoretical vs. Non-Theoretical approaches. The main difference between the theoretical and non-theoretical approaches in this context is - Theoretical approaches use TDA (sometimes in combination with other numerical techniques) to understand and explain linguistic phenomena, while Non-theoretical approaches use TDA to enhance or explain model performance. However, linguistic phenomena understood from the Theoretical approaches can inform Non-theoretical approaches.
@@ 3.1.1 Semantic Topological Space
In addition, Cavaliere et al. (2017) extracts main concepts from the texts by probing the context-aware semantic topological space built with simplicial complexes. Finally, Wagner et al. (2012) uses TF-IDF to numerically represent the top 10-50 words in a corpus and build a topological space that analyzes the structure of similarities within several documents. This topological space is built using discrete Morse theory and persistent homology to find meaningful topological patterns Wagner et al. (2012).
Insight (Semantic Space). The semantic topological space is explored by researchers to identify semantic linguistic principles captured in texts through a topological lens. Most of the applications in this section involve understanding the semantic similarity between text pairs from a topological lens.
@@ 3.1.2 Syntactic Topological Space
Therefore, Port et al. (2018) analyzes how syntactic parameters are distributed over different language families, including Indo-European, Niger-Congo, Austronesian, and Afro-Asiatic families. , and features in $\beta_{1}$ capture syntactic differences between branches of families of languages, as well as the syntactic influences between them Port et al. (2018). Port et al. (2018) show that the three persistent connected components ($\beta_{0}$) in the Niger-Congo family represents its three subfamilies - Mande, Atlantic-Congo, and Kordofania. The syntactic topological structures of these languages also reveal the historical linguistic phenomena that the Hellenic branch played a role in the historical development of the Indo-European languages Port et al. (2018).
@@
Insight (Syntactic Space). The syntactic topological space captures the syntactic structure of language (i.e., grammar, etc.) from a topological lens. Using this framework, researchers confirm linguistic phenomena in language families and subfamilies by exploring the syntactic relationship between languages. Thus, such a framework can be further utilized to discover novel linguistic phenomena.
@@ 3.2 Topology of Topic Evolution
The topology of topic evolution refers to the study and representation of how topics, themes, or concepts develop and change over time within a given corpus of texts or discourses in a topological space/structure. This concept is particularly relevant in fields where understanding the temporal dynamics of topics can provide insights into trends, shifts in public opinion, or the development of scientific or cultural themes.
Thus, Sami and Farrahi (2017) utilizes TDA to visualize the relationship between words in a text block, words in a corpus, and text blocks in a corpus. They visualize both local context (i.e., each text block in a set of sentences), and global context (i.e., occurrence of extracted words in the corpus) features. These features are extracted by using the circular topology to represent words. Finally, findings reveal that using the circular topology in 2D space, core words from the corpus stay close to the center, and the explanatory words remain close to the circle’s periphery.
Insight (Topic Evolution). Exploring the topology of topic evolution is a novel framework for capturing the topology of topics in a corpus. The findings suggest that this framework can be adopted to evaluate the utility of a summarization, paraphrasing, or obfuscating model, by comparing the topology of the topic evolution of the original vs. the perturbed texts.
@@ 3.3 Topological “Shape” of Words
Draganov and Skiena (2024) captures the “shape” of words for several languages by comparing the phylogenies or evolutionary history of language in the Indo-European language family. Initially, numerically representing the texts with FastText, they use persistent homology to construct language phylogenetic trees for over 81 Indo-European languages. Experiments reveal that: (1) the shape of the word embedding of a language carries historical and structural information, similar to Port et al. (2018, 2022)’s findings; and (2) TDA methods can successfully capture aspects of the shape of language embeddings Draganov and Skiena (2024). Next, Fitz (2022) introduces a novel terminology - word manifold, which is a simplicial complex, whose topological space captures grammatical structure expressed by the corpus. This is done by implementing a technique for generating topological structure directly from strings of words. Experiments reveal that the homotopy type of the word manifold is also influenced by linguistic structure Fitz (2022). Similarly, Fitz et al. (2024) measures the topological complexity (known as perforation) of hidden representation of Transformer-based models to understand their topological shapes.
Similar to Port et al. (2018, 2022), Dong (2024) extracts the topological shapes of languages. Specifically South American languages - the Nuclear-Macro-Jê (NMJ) and Quechuan families using TDA. By using techniques like multiple correspondence analysis (MCA) for dimension reduction of the categorical-valued dataset and persistent homology, Dong (2024) visualizes each language in the selected families as a point cloud. This forms the topological shape of the South American languages, such that languages close together are more similar. By comparing the topological shapes of the languages, it is observed that there are major distinctions between the Jê-proper and the non-Jê-proper languages, as well as the northern and southern Quechuan languages Dong (2024). Finally, Bouazzaoui et al. (2021) explores the topological similarity of the shapes of two writing systems - Tifinagh and Phoenician scripts.
Insight (Shape of Words). The topological “shape” of words is a concept that has interested several linguists, as it can be used to confirm and discover linguistic phenomena within languages. It is the most studied theoretical approach, with the main application focused on capturing the shape of several languages. This concept combines all other frameworks like the semantic and syntactic topological spaces to capture a linguistically informed topological shape of texts.
@@ 4.1 Tasks, Learning Types, Modality, and TDA techniques
We categorize the NLP tasks to which TDA has been applied into seven categories:
  1. 1.
    Classification: The most popular application is deepfake text detection Løvlie (2023); Tulchinskii et al. (2024); Kushnareva et al. (2024, 2021); Uchendu et al. (2024); Wei et al. (2025).
  2. 2.
    Clustering and Topic Modeling: The most popular application is document clustering and topic modeling Holmes (2020); Guan et al. (2016).
  3. 3.
    Sentiment and Semantic Analysis: The most popular applications are linguistic/grammatical acceptability Cherniavskii et al. (2022); Jain et al. (2024), word sense induction and disambiguation Rawson et al. (2022); Temčinas (2018), and polysemy word classification Jakubowski et al. (2020); Shehu (2024).
  4. 4.
    Structure and Visualization: The most popular is using Mapper to visualize model hidden weights Garcia (2022); Rathore et al. (2023).
  5. 5.
    Health, Social, and Scholarly Analysis: Since this is not a popular application for TDA, the most interesting applications are - prediction of epidemics Petri and Leitao (2020) and categorization of lonely people Effah (2017).
  6. 6.
    Speech and Music Processing: The most popular applications are studying vocalizations Bonafos et al. (2023, 2024) and music classification Bergomi (2015); Sassone et al. (2022).
  7. 7.
    Model Interpretation and Analysis: The most popular applications are model probing to reveal behavior in hidden weights Kostenok et al. (2023); Gourgoulias et al. (2024).
Categories (Tasks). The seven categories of Tasks are selected based on surveying the types of problems researchers have attempted to solve by employing TDA techniques.
Finally, we can also categorize non-theoretical TDA applications in NLP by learning types, Modality, and TDA techniques. Learning types has supervised Elyasi and Moghadam (2019); Lavery et al. (2024), and unsupervised Spannaus et al. (2024); Bonafos et al. (2024); Modality has text Triki (2021); Kostenok et al. (2023), and Speech Sassone et al. (2022); and TDA techniques, has Persistent Homology Torres-Tramón et al. (2015); Cherniavskii et al. (2022), and Mapper Holmes (2020); Elyasi and Moghadam (2019).
Categories (Learning Types, Modality, & TDA techniques). Due to the binary nature of these three categories, they are not very informative and thus unsuitable as the main taxonomy for the non-theoretical applications.